OpenStudy (goodvibes101):

What is the kinetic energy of a 5 kg fish swimming at 6 km/h? *Please show work*

OpenStudy (peachpi):

\[K=\frac{ 1 }{ 2 }mv^2\] Probably ought to convert 6 km/h to m/s

OpenStudy (goodvibes101):

Thank you . So K=1/2 mv^2 Is the answer ?

OpenStudy (peachpi):

no that's the formula. You need to plug the number in. m = 5 kg For v, you have to convert 6 km/h to m/s before using it in the formula.

OpenStudy (goodvibes101):

Lol. So what's the answer? Im in a bit of a rush .

OpenStudy (goodvibes101):

@peachpi ..

OpenStudy (anthonyym):

In the fromula\[K=\frac{ 1 }{ 2 }mv^2\], K stands for kinetic energy, m stand for mass, and v stands for velocity. From your given information, the mass is 5 kg, and the velocity is 6 km/h, or 5/3 m/s

OpenStudy (goodvibes101):

So the answer would be : K= 30 kg ? Or Am I doing something wrong ? @anthonyym

OpenStudy (mjdennis):

30 is not the correct value, and kg isn't a unit of energy. Do you still want help on this?

OpenStudy (shamim):

k=?

OpenStudy (shamim):

given mass m=10kg

OpenStudy (shamim):

velocity v=6km/h=?m/s

OpenStudy (shamim):

u know 1 km=1000m

OpenStudy (shamim):

1 h=3600sec

OpenStudy (shamim):

so v=6*1000m/3600sec=?

OpenStudy (shamim):

u hv to try to understand the fact

OpenStudy (shamim):

not try to get the result ony

OpenStudy (reemj):

K = 1/2 m v^2 K = 1/2 ( 5 ) ( 6 )^2 K = 1/2 ( 5 ) ( 36 ) K = 90

OpenStudy (mjdennis):

@Reemj: Your answer is NOT correct. You can't just plug the numbers without the units. PLEASE READ AND UNDERSTAND the content of other people's responses; the responses above contain corrections for the difference between km/hr and m/s. PLEASE READ the Code of Conduct, and try to help the questioner learn something, instead of handing out an answer that gives no clarity. PLEASE CONSIDER the context -- the question has been ignored by the questioner for 5 days, after much insistence that he/she was in a rush.